OpenStudy (anonymous):

Evaluate the following limits. If needed, enter 'infinity' for \infty and '-infinity' for -\infty. \lim_{ x \rightarrow \infty } \frac {\sqrt{ 10 + 9 x^2 } }{ 5 + 7 x } = \lim_{ x \rightarrow -\infty } \frac {\sqrt{ 10 + 9 x^2 } }{ 5 + 7 x } =

OpenStudy (anonymous):

\[\lim_{ x \rightarrow \infty } \frac {\sqrt{ 10 + 9 x^2 } }{ 5 + 7 x } =\]

OpenStudy (anonymous):

isnt it the first one is infinity? but the answer is wrong

OpenStudy (anonymous):

\[\frac{\sqrt{9x^2}}{7x}=\frac{3x}{7x}=\frac{3}{7}\] is the eyeball way

OpenStudy (anonymous):

Oh!!! just take the greatest power?!

OpenStudy (anonymous):

yeah the other stuff is not important if you have a ratio and \(x\to \infty\) you can do it a longer more involved way to make a math teacher happy, but there is no need for it

OpenStudy (anonymous):

\[\lim_{ x \rightarrow -\infty } \frac {\sqrt{ 10 + 9 x^2 } }{ 5 + 7 x } = \] you have to be a bit more careful, because \(x\to -\infty\) and so \(7x<0\) whereas \(\sqrt{9x^2}>0\) so the answer is not \(\frac{3}{7}\) but rather \(-\frac{3}{7}\)

OpenStudy (anonymous):

Oh yea so the answer for the second one is -3/7

OpenStudy (anonymous):

Yes!! thx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

what about \lim_{ x \to \infty } \frac{ 3 }{ e^x + 8 } =

OpenStudy (anonymous):

@satellite73 I know that if the limit go with -infty is will be 3/8 but how about the positive infty

OpenStudy (anonymous):

\[\lim_{ x \to \infty } \frac{ 3 }{ e^x + 8 }\]

OpenStudy (anonymous):

\(e^x\) goes to infinity lickety split, so that limit is zero

OpenStudy (anonymous):

and you are right with the first one, since \(\lim_{x\to -\infty}e^x=0\)

OpenStudy (anonymous):

Yea!!!!!! oh thx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

but im still confuse y limx→−∞e^x=0

OpenStudy (anonymous):

b/c e^x is always 0?

OpenStudy (anonymous):

think about \(e^{-100}=\frac{1}{e^{100}}\) it is an amazingly small number (i.e. close to zero) since \(e^100\) is really really large

OpenStudy (anonymous):

oh yea it is close to 0 kk i got it REALLY thanks!!!